OpenStudy (anonymous):

completing the square solve 2x^2+x=3

OpenStudy (anonymous):

2x^2+x-3=0 (2x+3)(x-1)=0 x= -3/2 or 1

OpenStudy (anonymous):

thats factoring... not completing the square. :(

OpenStudy (anonymous):

ahh sorry didnt read properly

OpenStudy (anonymous):

i know how to factor it out i just totally forgot how to complete the square.

OpenStudy (anonymous):

ok lemme do again

OpenStudy (anonymous):

alright. kk

OpenStudy (anonymous):

2x^2+x-3=0 x^2+x/2-3/2=0 x^2+x/2+(x^2/16)=3/2+1/16 (x+1/4)^2=25/16

OpenStudy (anonymous):

did you had to square root?

OpenStudy (anonymous):

and then simply rearrange to get x=plus/minus 5/4 -1/4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i skipped that part cuz i was slow at typing and didnt want u to die or boredorm

OpenStudy (anonymous):

of*

OpenStudy (anonymous):

could u draw it?

OpenStudy (anonymous):

haha ok ill try

OpenStudy (anonymous):

sorry i learn from seeing a picture

OpenStudy (anonymous):

please correct me if im wrong in any areas

OpenStudy (anonymous):

nah its all goods. Most of us are visual learners =D

OpenStudy (anonymous):